Sheridan Enterprises sells a product for $80 per unit. The variable cost is $43 per unit, while fixed costs are $235,468.

Determine (a) the break-even point in sales units and (b) the break-even point if the selling price were increased to $86 per unit.

a. Break-even point in sales units units
b. Break-even point if the selling price were increased to $86 per unit units
Answer #1

(a)

Sales price per unit = $ 80
Variable cost per unit= $ 43
Contribution per unit = $ 37

Fixed costs = $ 235,468

Breakeven point = Fixed costs / Contribution per unit = 235,468 / 37 = 6,364 units

(b)

Sales price per unit = $ 86
Variable cost per unit = $ 43
Contribution per unit = $ 43

Breakeven point = 235,468 / 43 = 5,476 units